OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at \((0,−9)\) and a directrix \(y=9.\)

OpenStudy (welshfella):

this parabola opens upwards and standard form is x^2 = 4ay where a is the y coordinate of the focus and y = -a is the directrix so can you work this one out?

OpenStudy (anonymous):

Wouldn't the parabola open downwards?

OpenStudy (anonymous):

It would. The form for a parabola is \[y = 1/4p(x-h)^2+k\] From here you can plug 9 into p since p = distance between vertex and focus. The h and k values are both 0 since the vertex is at (0,0). Can you get it from here?

OpenStudy (anonymous):

I lied, plug -9 in for p.

OpenStudy (anonymous):



OpenStudy (anonymous):

P is (x,y)

OpenStudy (anonymous):

Hang on, I'm trying to solve it. :)

OpenStudy (anonymous):

Okay, post your answer here so I can check it for you :)

OpenStudy (anonymous):

So is it \(y=-\frac{ 1 }{ 36 }x^2?\)

OpenStudy (anonymous):

It sure is!